### v1

@@ 4 Conclusion
We also showed that a sufficient winning condition for the second player for Arc-Kayles is GI-hard, which is one of the few hardness results on this game. While this is not enough to determine the computational complexity of Arc-Kayles, such an approach might help to understand more about it. A future research direction would be to study more necessary and sufficient conditions for first and second player winning, and determine their complexity.
On the polynomial-time side, we showed that Non-Disconnecting Arc-Kayles is easy on tree-like graphs, such as unicyclic graphs and clique trees. Note that this contrasts with Arc-Kayles, where even subdivided stars with three paths are open. Future research could consider other tree-like classes, such as cacti. Furthermore, our study of clique-twin-free threshold graphs seems to imply that the complexity gap may lie between threshold and split graphs. Another direction would be to look at cographs, since threshold graphs are the intersection of split (for which the game is PSPACE-complete) and cographs.

### v2

@@ 4 Conclusion
We also showed that a natural sufficient winning condition for the second player for Arc-Kayles is GI-hard, which is one of the few hardness results on this game. While this is not enough to determine the computational complexity of Arc-Kayles, such an approach might help to understand more about it. A future research direction would be to study more necessary and sufficient conditions for first and second player winning, and determine their complexity.
On the polynomial-time side, we showed that Non-Disconnecting Arc-Kayles is easy on tree-like graphs, such as clique trees, and gave a polynomial kernel parameterized by feedback edge number. Note that this contrasts with Arc-Kayles, where even subdivided stars with three paths are open. Future research could consider other tree-like classes with unbounded feedback edge number, such as cacti. Furthermore, our study of clique-twin-free threshold graphs seems to imply that the complexity gap may lie between threshold and split graphs. Another direction would be to look at cographs, since threshold graphs are the intersection of split (for which the game is PSPACE-complete) and cographs.